Calculate the monthly finance charge for the credit card transaction. Assume that it takes 10 days for a payment to be received and recorded, and that the month is 30 days long. (Round your answer to the nearest cent.)

$350 balance, 17%, $300 payment; average daily balance method
Answer #1

Average daily balance Method

Number of days for which $350 Balance is Held = 10 Days

Now, The Balance remaining on 10th of month = ( 350 - 300)

= 50

Number of days for which $50 Balance is Held = 20 Days

So, Average balance = ( 350 * 10 + 50*20) / 30

= 150

Finance charge = Average balance * ( Interest rate per month)

= 150 * (17%/12)

Finance charge = $2.125